# Green Federated Learning: A new era of Green Aware AI

### v1

@@ 5.1. Non-IID Data Handling FL
A new distributed machine learning platform for protecting privacy is called federated learning. However, when the training data are not independently and identically distributed (Non-IID) on the local devices, models built in FL typically perform worse than those trained in the conventional centralized learning mode (Zhu et al., 2021). To handle this issue, several authors have proposed different FL approaches. In (McMahan et al., 2016), the authors proposed a simple FL algorithm, FedAvg, to handle non-IID data. The authors’ comprehensive testing of this method demonstrates that it can significantly reduce the number of communication rounds needed to train a deep network using decentralized data. It is resilient to non-IID and imbalanced data distributions. FedAvg trains high-quality models with very few communication rounds, as demonstrated by results on a variety of model architectures, such as a multi-layer perceptron, two different convolutional NNs, a two-layer character LSTM, and a large-scale word-level LSTM. However, the authors in (Sahu et al., 2018) claimed that the FedAvg failed to give the convergence guarantee with non-IID data and addressed heterogeneity in federated networks by introducing the FedProx framework. However, FedProx can be seen as a generalization and re-parametrization of FedAvg. Although the approach is only slightly altered by this re-parameterization, the changes have significant implications for theory and practice. FedProx is guaranteed to converge theoretically when learning over data from non-identical distributions (statistical heterogeneity) and when complying with device-level systems constraints by letting each participating device do a different amount of work (systems heterogeneity). In practice, on various real-world federated datasets, FedProx enables more robust convergence than FedAvg. Specifically, FedProx shows far more accurate and persistent convergence behavior than FedAvg, enhancing absolute test accuracy by 22% on average in extremely heterogeneous situations. At the same time, in (Zhao et al., 2018), the authors experimentally demonstrated that when neural networks are trained with highly skewed non-IID data, where each client device trains just on a single class of data, the accuracy of federated learning (FedAvg) decreases significantly by as much as  55%. To address this, the authors suggested FedCurv (Shoham et al., 2019), an FL-based method that enhances training on non-IID data: developing a limited subset of data that is globally shared among all edge devices. In (Karimireddy et al., 2020), the authors demonstrated that when the data is heterogeneous (non-iid), FedAvg experiences “client-drift,” which leads to unstable and sluggish convergence. In order to address this, the authors in (Karimireddy et al., 2020) suggested a brand-new technique called SCAFFOLD, which makes use of control variates (variance reduction) to adjust for “client-drift” in local updates. SCAFFOLD is insensitive to client sampling and data heterogeneity, requiring a notably smaller number of communication rounds. Additionally, SCAFFOLD can benefit from data similarity in the client, resulting in even faster convergence for quadratic. FedPer (Arivazhagan et al., 2019b) is a personalized FL framework which also handles the non-i.i.d. data in an efficient manner.
@@
FedAvg (McMahan et al., 2016) X X X X X
FedProx (Sahu et al., 2018) X X X X
FedPer (Arivazhagan et al., 2019b) X X X X
FedCurv (Shoham et al., 2019) X X X
SCAFFOLD(Karimireddy et al., 2020) X X
MOON (Li et al., 2021a) X X X
FedNova (Wang et al., 2020a) X X X
FedPAQ (Reisizadeh et al., 2019) X X X
FEDEMD (Huang et al., 2023a) X X
FedDyn (Acar et al., 2021) X X X
FedMA (Wang et al., 2020b) X X X
FedPD (Zhang et al., 2020) X X
5.2. Convergence
@@
FedNova (Wang et al., 2020a) is a quick error convergence normalized averaging method that removes objective inconsistency. This method mainly focuses on the heterogeneity of local updates of the model. It gives the first systematic understanding of the solution bias and the convergence delay caused by objective inconsistency, and it incorporates previously presented methods like FedAvg and FedProx. SCAFFOLD (Karimireddy et al., 2020) is another proposed FL algorithm that handles both the non-IID data and convergence rate. The authors in (Karimireddy et al., 2020) proved that the FedAvg FL algorithm suffers from client drift while using heterogeneous (non-i.i.d) data with stochastic Gradient Descent (SGD) as a local solver, resulting in slow convergence. The authors offered SCAFFOLD as a remedy that accounts for the client drift in its local updates using control variates (variance reduction). Moreover, SCAFFOLD is insensitive to client sampling and data heterogeneity, requiring a notably smaller number of communication rounds. Model-contrastive learning (MOON) (Li et al., 2021a) is another FL method that can handle non-i.i.d data and convergence rate simultaneously. MOON is a unique approach to addressing the non-IID data issue and a straightforward and efficient FL framework.
5.3. Communication-efficiency
@@
The authors in (Reisizadeh et al., 2019) proposed a communication-efficient FL method known as FedPAQ. FedPAQ is based on periodic averaging and quantization to overcome the communication bottleneck and scalability challenges. FedPAQ is based on three main features: (1) partial device participation, where only a portion of the devices take part in each training round; (2) periodic averaging, where models are updated locally at devices and only periodically averaged at the server; and (3) quantized message passing, where the edge nodes quantize their updates before uploading to the parameter server. These aspects address the scalability and communication issues in federated learning. We additionally illustrate the near-optimal theoretical guarantees that FedPAQ achieves for both strongly convex and non-convex loss functions and the empirical tradeoff between computation and communication. Federated Matched Averaging (FedMA) (Wang et al., 2020b) is another FL approach that reduces communication rounds efficiently with heterogeneous data. For contemporary CNNs and LSTMs, FedMA is a novel layers-wise federated learning approach that leverages Bayesian nonparametric techniques to adjust to data heterogeneity. The authors provided empirical evidence demonstrating that FedMA outperforms the most advanced FL algorithms while also reducing the overhead of communications. FedDyn (Acar et al., 2021), a dynamic regularization FL technique. A novel idea at the core of FedDyn is the dynamic updating of each device’s risk objective in each round to guarantee that the device optima are asymptotically compatible with the stationary points of the global empirical loss. The authors obtained sharp findings for communication rounds necessary to achieve target accuracy, and they proved to have convergence results for FedDyn in both convex and non-convex scenarios. The convex case results significantly outperform previous state-of-the-art efforts. In benchmark examples, FedDyn consistently achieves considerable communication savings over competing techniques on massively distributed large-scale text data, spanning different options of device participation and heterogeneity. The quality of communication lines, heterogeneity, and massively distributed data do not theoretically affect FedDyn. Federated Primal-Dual (FedPD) (Zhang et al., 2020) attains convergence under specific assumptions and the best possible optimization and communication complexity when the data is not i.i.d.
@@
FedCS (Nishio and Yonetani, 2018) X Before a deadline, select as many clients as you can. It is predicated on a knapsack-constrained greedy method.
@@
(Rizk et al., 2021) X Two-tier significance sampling for both data and clients. Clients are chosen initially, and then their data is sampled.
@@
OCEAN (Xu and Wang, 2021) X Allocation of bandwidth based on energy limits of the client. In order to improve the client selection pattern, it makes use of wireless channel information.
@@
Oort(Lai et al., 2021a) Make use of client-side heterogeneous systems and data. In order to choose participants based on their resilience to outliers, it uses an exploration-exploitation technique.
@@
(Amiri et al., 2021) Learn about local gradient norm and shared block fading wireless channels. It creates a policy for allocating resources in order to plan low-profile client devices.
@@
FOLB (Nguyen et al., 2021) Based on how well the local and global updates correlate. It makes an estimate of the clients’ potential to contribute to the updates.
@@
(Chen et al., 2022) X To address the communication bottleneck, employ norm-based client selection. It uses a basic algorithm to approximate the ideal client selection formula.
@@
FEDPNS(Wu and Wang, 2022) X Removing negative local updates through a comparison of the global and local gradients. Clients that accelerate model convergence are given preference.
@@
Power-of-Choice (Jee Cho et al., 2022) Balance between solution bias and convergence speed in local loss-based client selection. It demonstrates how skewed client selection might hasten convergence.
@@
PyramidFL (Li et al., 2022) Take advantage of system and data heterogeneity within a chosen clientele. It then selects the clients depending on their utilities and optimizes utility profiling locally.
@@
FCFL (Zhou et al., 2022) Wearable technology under unfavorable networking circumstances. It suggests aggregating only the model updates with the highest contributions using movement-aware FL.
@@
Eiffel (Sultana et al., 2022) Together, take into account elements like the age of the update and resource requirements. The frequency of global and local model updates is adaptively adjusted.
@@
(Luo et al., 2022) Customized client sampling probabilities are optimized to handle statistical and system heterogeneity. The wall-clock convergence time is reduced.
@@
F3AST (Ribero et al., 2023) X Acquires knowledge of a client selection method that is contingent on availability in order to reduce the effect of client-selection variance on the convergence of the global model.
@@
(Ma et al., 2023) X Provide a block-wise incremental gradient aggregation approach along with a dynamic user and task scheduling scheme.
5.4. Client-selection
@@
The FL process’s crucial step of client selection may impact the convergence rate, model precision, and clients’ overall energy usage. Moreover, the heterogeneity problem can be tackled by an efficient client selection algorithm (Fu et al., 2023b). A successful FL client selection technique can strengthen robustness (Nguyen et al., 2021), increase fairness (Sultana et al., 2022), decrease training overheads (Li et al., 2022), and improve model accuracy (Lai et al., 2021b). As a result, FL client selection research has been developing quickly in the last few years, according to the research community (Soltani et al., 2022). Many strategies seek to choose the most significant number of clients who can complete their model training and upload before the cutoff date in each global iteration. Federated learning needs to choose the right clients at the right time in order to speed up convergence and increase accuracy. In (Huang et al., 2023a), the authors provided FEDEMD, an adaptive client selection method based on the local and global data distributions’ Wasserstein distance. When there are multiple types of Mavericks present, FEDEMD reliably assures quick convergence by adjusting the selection probability so that Mavericks are preferred when the model gains from improvement on rare classes. When compared to current methods, such as those based on Shapley values, FEDEMD enhances neural network classifier convergence by at least 26.9% for FedAvg aggregation when compared to the state of the art. All the client selection strategies mentioned in Table 4 maintain the trade-off between communication and computation efficiency. However, reduced communication rounds and computation efficiency minimize the energy consumption of an FL model. Hence, make a green FL model.
5.5. Model Compression
@@
Model compression is proposed to minimize the neural network’s complexity before it is deployed for edge devices with restricted resources. In 1988, the first studies on model pruning were conducted. An amplitude-based pruning technique for shallow, fully linked networks was presented by Hanson and Pratt (Hanson and Pratt, 1988). Recently, in order to compress CNNs, the authors in (Han et al., 2015a) used many techniques, including quantization, pruning, and Huffman coding. Summing the convolution kernel’s absolute value was suggested by Li et al. (Li et al., 2017) as a criterion for determining its significance. This was the first instance of modifying the model structure to accomplish network compression using the convolution kernel as the pruning unit. Several criteria are proposed by researchers in (Lin et al., 2020; Li et al., 2017; Liu et al., 2015, 2017; He et al., 2017; Molchanov et al., 2017; Lee et al., 2019; He et al., 2019; Guo et al., 2020) and others to assess the role of convolution kernels in structured pruning; superfluous convolution kernels are eliminated or set to zero. To find redundant parameters during training, three pruning techniques—iterative pruning (Han et al., 2015b), gentle pruning (He et al., 2020), and dynamic pruning (Shen et al., 2021)—were suggested. Currently, there is no consensus on assessing the model’s parameter efficacy while minimizing accuracy loss.
@@
FL-PQSU (Xu et al., 2021) X It is made up of a three-stage pipeline that combines weight quantization, structured pruning, and selective updating to lower computing, storage, and communication costs and speed up FL training.
@@
FedTiny (Huang et al., 2023b) X X Produced customized small models for participating devices with sensitive local data that are limited in terms of memory and computation. Presented an adaptive batch normalization (BN) selection module to adaptively get an initially trimmed model to meet deployment circumstances, mitigating the effects of biased pruning caused by unseen heterogeneity data over devices.
@@
(Shah and Lau, 2023) X X In order to tackle, an infeasible volume of data, compression methods to build sparse models from dense networks with much less bandwidth and storage needed. To do this, compression methods for the client models, which deal with upstream communication, and the server model, which deals with downstream communication have been considered. These two are essential for creating and preserving sparsity over communication cycles.
@@
FedCOMGATE (Haddadpour et al., 2020) X It is a federated averaging variation for heterogeneous environments that includes compression and local gradient tracking. For generic non-convex, strongly convex, and convex objectives, it can determine its convergence rates.
@@
FedHBAA—FL (Chen et al., 2024a) X X combined server aggregation technique with quantization bit allocation. A server aggregation technique in FedHBAA for any given client bit allotment. For both the strongly convex and nonconvex loss functions, a closed-form aggregation weight solution as a function of the allotted bits is achieved by minimizing the drift term in the convergence rate analysis. The best bit allocation plan is then found by solving the derived optimal aggregation weights.
@@
AGQFL (Lian et al., 2021) X X Three components make up this automatic gradient quantization method: the quantization indicator, quantization strategy, and quantization optimizer. By evaluating the current model’s capacity for convergence, the quantization indicator module automatically ascertains the direction in which the quantization precision should be adjusted. The quantization strategy module modifies the quantization precision at run-time based on the indication and each node’s physical bandwidth. Additionally, in order to minimize training bias and get rid of instability during the training phase, the quantization optimizer module creates a new optimizer.
@@
NQFL (Chen et al., 2024b) X X The significant communication overhead that comes with FL is still a major obstacle. This letter introduces a nonuniform quantization approach based on the Lloyd-Max algorithm to address this problem. This method uses fewer communication resources to accomplish the same result as other methods. Through numerical simulations and performance analysis, we confirm the convergence and efficiency of the suggested approach. It illustrates how our method can save communication overhead while preserving dependable FL system performance.
@@
FedQNN (Ji and Chen, 2023) X A communication-and computation-efficient FL framework for Internet of Things applications. For the first time, ultralow-bitwidth quantization is integrated into the FL environment, enabling clients to carry out lightweight fix-point computing effectively and with reduced power consumption. Furthermore, by combining quantization and sparsification techniques, both upstream and downstream data are greatly reduced for more effective communication.
@@
FedDQ (Qu et al., 2022) X X A descending quantization system, is proposed based on the theoretical analysis that the range of the model updates will progressively contract because the model is meant to converge with training, meaning that the quantization level should decrease as training progresses.
@@
AdaQuantFL (Jhunjhunwala et al., 2021) X X An adaptive quantization technique modifies the number of quantization levels throughout training with the goal of achieving both a low error floor and effective communication.
@@
FedQOGD (Park et al., 2022) X X Gradient descent quantized communication efficient FL framework using partial node participation and stochastic quantization. FedQOGD has a minimal communication overhead and maintains edge-node privacy while yielding the same asymptotic performance as its centralized equivalent (i.e., all local data are gathered at the central server).
@@
In order to alleviate the communication delay constraint in federated learning, gradient compression—which includes gradient quantization and sparsification—has been extensively studied. Quantization modifies a floating-point number of bits to compress parameters (Tonellotto et al., 2021). Low-precision representation of deep neural networks (DNNs) is a potential method to allow for effective memory reduction and acceleration. Hence, the reduction in energy consumption.
6. Energy and Carbon Footprint Study in Federated Learning

### v2

@@ Introduction
A new distributed machine learning platform for protecting privacy is called federated learning. However, when the training data are not independently and identically distributed (Non-IID) on the local devices, models built in FL typically perform worse than those trained in the conventional centralized learning mode [Zhu et al.(2021)]. To handle this issue, several authors have proposed different FL approaches. In [McMahan et al.(2016)], the authors proposed a simple FL algorithm, FedAvg, to handle non-IID data. The authors’ comprehensive testing of this method demonstrates that it can significantly reduce the number of communication rounds needed to train a deep network using decentralized data. It is resilient to non-IID and imbalanced data distributions. FedAvg trains high-quality models with very few communication rounds, as demonstrated by results on a variety of model architectures, such as a multi-layer perceptron, two different convolutional NNs, a two-layer character LSTM, and a large-scale word-level LSTM. However, the authors in [Sahu et al.(2018)] claimed that the FedAvg failed to give the convergence guarantee with non-IID data and addressed heterogeneity in federated networks by introducing the FedProx framework. However, FedProx can be seen as a generalization and re-parametrization of FedAvg. Although the approach is only slightly altered by this re-parameterization, the changes have significant implications for theory and practice. FedProx is guaranteed to converge theoretically when learning over data from non-identical distributions (statistical heterogeneity) and when complying with device-level systems constraints by letting each participating device do a different amount of work (systems heterogeneity). In practice, on various real-world federated datasets, FedProx enables more robust convergence than FedAvg. Specifically, FedProx shows far more accurate and persistent convergence behavior than FedAvg, enhancing absolute test accuracy by 22% on average in extremely heterogeneous situations. At the same time, in [Zhao et al.(2018)], the authors experimentally demonstrated that when neural networks are trained with highly skewed non-IID data, where each client device trains just on a single class of data, the accuracy of federated learning (FedAvg) decreases significantly by as much as  55%. To address this, the authors suggested FedCurv [Shoham et al.(2019)], an FL-based method that enhances training on non-IID data: developing a limited subset of data that is globally shared among all edge devices. In [Karimireddy et al.(2020)], the authors demonstrated that when the data is heterogeneous (non-iid), FedAvg experiences “client-drift,” which leads to unstable and sluggish convergence. In order to address this, the authors in [Karimireddy et al.(2020)] suggested a brand-new technique called SCAFFOLD, which makes use of control variates (variance reduction) to adjust for “client-drift” in local updates. SCAFFOLD is insensitive to client sampling and data heterogeneity, requiring a notably smaller number of communication rounds. Additionally, SCAFFOLD can benefit from data similarity in the client, resulting in even faster convergence for quadratic. FedPer [Arivazhagan et al.(2019b)] is a personalized FL framework which also handles the non-i.i.d. data in an efficient manner.
@@
FedAvg [McMahan et al.(2016)] X X X X X
FedProx [Sahu et al.(2018)] X X X X
FedPer [Arivazhagan et al.(2019b)] X X X X
FedCurv [Shoham et al.(2019)] X X X
SCAFFOLD[Karimireddy et al.(2020)] X X
MOON [Li et al.(2021a)] X X X
FedNova [Wang et al.(2020a)] X X X
FedPAQ [Reisizadeh et al.(2019)] X X X
FEDEMD [Huang et al.(2023a)] X X
FedDyn [Acar et al.(2021)] X X X
FedMA [Wang et al.(2020b)] X X X
FedPD [Zhang et al.(2020)] X X
Convergence
@@
FedNova [Wang et al.(2020a)] is a quick error convergence normalized averaging method that removes objective inconsistency. This method mainly focuses on the heterogeneity of local updates of the model. It gives the first systematic understanding of the solution bias and the convergence delay caused by objective inconsistency, and it incorporates previously presented methods like FedAvg and FedProx. SCAFFOLD [Karimireddy et al.(2020)] is another proposed FL algorithm that handles both the non-IID data and convergence rate. The authors in [Karimireddy et al.(2020)] proved that the FedAvg FL algorithm suffers from client drift while using heterogeneous (non-i.i.d) data with stochastic Gradient Descent (SGD) as a local solver, resulting in slow convergence. The authors offered SCAFFOLD as a remedy that accounts for the client drift in its local updates using control variates (variance reduction). Moreover, SCAFFOLD is insensitive to client sampling and data heterogeneity, requiring a notably smaller number of communication rounds. Model-contrastive learning (MOON) [Li et al.(2021a)] is another FL method that can handle non-i.i.d data and convergence rate simultaneously. MOON is a unique approach to addressing the non-IID data issue and a straightforward and efficient FL framework.
Communication-efficiency
@@
The authors in [Reisizadeh et al.(2019)] proposed a communication-efficient FL method known as FedPAQ. FedPAQ is based on periodic averaging and quantization to overcome the communication bottleneck and scalability challenges. FedPAQ is based on three main features: (1) partial device participation, where only a portion of the devices take part in each training round; (2) periodic averaging, where models are updated locally at devices and only periodically averaged at the server; and (3) quantized message passing, where the edge nodes quantize their updates before uploading to the parameter server. These aspects address the scalability and communication issues in federated learning. We additionally illustrate the near-optimal theoretical guarantees that FedPAQ achieves for both strongly convex and non-convex loss functions and the empirical tradeoff between computation and communication. Federated Matched Averaging (FedMA) [Wang et al.(2020b)] is another FL approach that reduces communication rounds efficiently with heterogeneous data. For contemporary CNNs and LSTMs, FedMA is a novel layers-wise federated learning approach that leverages Bayesian nonparametric techniques to adjust to data heterogeneity. The authors provided empirical evidence demonstrating that FedMA outperforms the most advanced FL algorithms while also reducing the overhead of communications. FedDyn [Acar et al.(2021)], a dynamic regularization FL technique. A novel idea at the core of FedDyn is the dynamic updating of each device’s risk objective in each round to guarantee that the device optima are asymptotically compatible with the stationary points of the global empirical loss. The authors obtained sharp findings for communication rounds necessary to achieve target accuracy, and they proved to have convergence results for FedDyn in both convex and non-convex scenarios. The convex case results significantly outperform previous state-of-the-art efforts. In benchmark examples, FedDyn consistently achieves considerable communication savings over competing techniques on massively distributed large-scale text data, spanning different options of device participation and heterogeneity. The quality of communication lines, heterogeneity, and massively distributed data do not theoretically affect FedDyn. Federated Primal-Dual (FedPD) [Zhang et al.(2020)] attains convergence under specific assumptions and the best possible optimization and communication complexity when the data is not i.i.d.
@@
FedCS [Nishio and Yonetani(2018)] X Before a deadline, select as many clients as you can. It is predicated on a knapsack-constrained greedy method.
@@
[Rizk et al.(2021)] X Two-tier significance sampling for both data and clients. Clients are chosen initially, and then their data is sampled.
@@
OCEAN [Xu and Wang(2021)] X Allocation of bandwidth based on energy limits of the client. In order to improve the client selection pattern, it makes use of wireless channel information.
@@
Oort[Lai et al.(2021a)] Make use of client-side heterogeneous systems and data. In order to choose participants based on their resilience to outliers, it uses an exploration-exploitation technique.
@@
[Amiri et al.(2021)] Learn about local gradient norm and shared block fading wireless channels. It creates a policy for allocating resources in order to plan low-profile client devices.
@@
FOLB [Nguyen et al.(2021)] Based on how well the local and global updates correlate. It makes an estimate of the clients’ potential to contribute to the updates.
@@
[Chen et al.(2022)] X To address the communication bottleneck, employ norm-based client selection. It uses a basic algorithm to approximate the ideal client selection formula.
@@
FEDPNS[Wu and Wang(2022)] X Removing negative local updates through a comparison of the global and local gradients. Clients that accelerate model convergence are given preference.
@@
Power-of-Choice [Jee Cho et al.(2022)] Balance between solution bias and convergence speed in local loss-based client selection. It demonstrates how skewed client selection might hasten convergence.
@@
PyramidFL [Li et al.(2022)] Take advantage of system and data heterogeneity within a chosen clientele. It then selects the clients depending on their utilities and optimizes utility profiling locally.
@@
FCFL [Zhou et al.(2022)] Wearable technology under unfavorable networking circumstances. It suggests aggregating only the model updates with the highest contributions using movement-aware FL.
@@
Eiffel [Sultana et al.(2022)] Together, take into account elements like the age of the update and resource requirements. The frequency of global and local model updates is adaptively adjusted.
@@
[Luo et al.(2022)] Customized client sampling probabilities are optimized to handle statistical and system heterogeneity. The wall-clock convergence time is reduced.
@@
F3AST [Ribero et al.(2023)] X Acquires knowledge of a client selection method that is contingent on availability in order to reduce the effect of client-selection variance on the convergence of the global model.
@@
[Ma et al.(2023)] X Provide a block-wise incremental gradient aggregation approach along with a dynamic user and task scheduling scheme.
Client-selection
@@
The FL process’s crucial step of client selection may impact the convergence rate, model precision, and clients’ overall energy usage. Moreover, the heterogeneity problem can be tackled by an efficient client selection algorithm [Fu et al.(2023b)]. A successful FL client selection technique can strengthen robustness [Nguyen et al.(2021)], increase fairness [Sultana et al.(2022)], decrease training overheads [Li et al.(2022)], and improve model accuracy [Lai et al.(2021b)]. As a result, FL client selection research has been developing quickly in the last few years, according to the research community [Soltani et al.(2022)]. Many strategies seek to choose the most significant number of clients who can complete their model training and upload before the cutoff date in each global iteration. Federated learning needs to choose the right clients at the right time in order to speed up convergence and increase accuracy. In [Huang et al.(2023a)], the authors provided FEDEMD, an adaptive client selection method based on the local and global data distributions’ Wasserstein distance. When there are multiple types of Mavericks present, FEDEMD reliably assures quick convergence by adjusting the selection probability so that Mavericks are preferred when the model gains from improvement on rare classes. When compared to current methods, such as those based on Shapley values, FEDEMD enhances neural network classifier convergence by at least 26.9% for FedAvg aggregation when compared to the state of the art. All the client selection strategies mentioned in Table 4 maintain the trade-off between communication and computation efficiency. However, reduced communication rounds and computation efficiency minimize the energy consumption of an FL model. Hence, make a green FL model.
Model Compression
@@
Model compression is proposed to minimize the neural network’s complexity before it is deployed for edge devices with restricted resources. In 1988, the first studies on model pruning were conducted. An amplitude-based pruning technique for shallow, fully linked networks was presented by Hanson and Pratt [Hanson and Pratt(1988)]. Recently, in order to compress CNNs, the authors in [Han et al.(2015a)] used many techniques, including quantization, pruning, and Huffman coding. Summing the convolution kernel’s absolute value was suggested by Li et al. [Li et al.(2017)] as a criterion for determining its significance. This was the first instance of modifying the model structure to accomplish network compression using the convolution kernel as the pruning unit. Several criteria are proposed by researchers in [Lin et al.(2020), Li et al.(2017), Liu et al.(2015), Liu et al.(2017), He et al.(2017), Molchanov et al.(2017), Lee et al.(2019), He et al.(2019), Guo et al.(2020)] and others to assess the role of convolution kernels in structured pruning; superfluous convolution kernels are eliminated or set to zero. To find redundant parameters during training, three pruning techniques—iterative pruning [Han et al.(2015b)], gentle pruning [He et al.(2020)], and dynamic pruning [Shen et al.(2021)]—were suggested. Currently, there is no consensus on assessing the model’s parameter efficacy while minimizing accuracy loss.
@@
FL-PQSU [Xu et al.(2021)] X It is made up of a three-stage pipeline that combines weight quantization, structured pruning, and selective updating to lower computing, storage, and communication costs and speed up FL training.
@@
FedTiny [Huang et al.(2023b)] X X Produced customized small models for participating devices with sensitive local data that are limited in terms of memory and computation. Presented an adaptive batch normalization (BN) selection module to adaptively get an initially trimmed model to meet deployment circumstances, mitigating the effects of biased pruning caused by unseen heterogeneity data over devices.
@@
[Shah and Lau(2023)] X X In order to tackle, an infeasible volume of data, compression methods to build sparse models from dense networks with much less bandwidth and storage needed. To do this, compression methods for the client models, which deal with upstream communication, and the server model, which deals with downstream communication have been considered. These two are essential for creating and preserving sparsity over communication cycles.
@@
FedCOMGATE [Haddadpour et al.(2020)] X It is a federated averaging variation for heterogeneous environments that includes compression and local gradient tracking. For generic non-convex, strongly convex, and convex objectives, it can determine its convergence rates.
@@
FedHBAA—FL [Chen et al.(2024a)] X X combined server aggregation technique with quantization bit allocation. A server aggregation technique in FedHBAA for any given client bit allotment. For both the strongly convex and nonconvex loss functions, a closed-form aggregation weight solution as a function of the allotted bits is achieved by minimizing the drift term in the convergence rate analysis. The best bit allocation plan is then found by solving the derived optimal aggregation weights.
@@
AGQFL [Lian et al.(2021)] X X Three components make up this automatic gradient quantization method: the quantization indicator, quantization strategy, and quantization optimizer. By evaluating the current model’s capacity for convergence, the quantization indicator module automatically ascertains the direction in which the quantization precision should be adjusted. The quantization strategy module modifies the quantization precision at run-time based on the indication and each node’s physical bandwidth. Additionally, in order to minimize training bias and get rid of instability during the training phase, the quantization optimizer module creates a new optimizer.
@@
NQFL [Chen et al.(2024b)] X X The significant communication overhead that comes with FL is still a major obstacle. This letter introduces a nonuniform quantization approach based on the Lloyd-Max algorithm to address this problem. This method uses fewer communication resources to accomplish the same result as other methods. Through numerical simulations and performance analysis, we confirm the convergence and efficiency of the suggested approach. It illustrates how our method can save communication overhead while preserving dependable FL system performance.
@@
FedQNN [Ji and Chen(2023)] X A communication-and computation-efficient FL framework for Internet of Things applications. For the first time, ultralow-bitwidth quantization is integrated into the FL environment, enabling clients to carry out lightweight fix-point computing effectively and with reduced power consumption. Furthermore, by combining quantization and sparsification techniques, both upstream and downstream data are greatly reduced for more effective communication.
@@
FedDQ [Qu et al.(2022)] X X A descending quantization system, is proposed based on the theoretical analysis that the range of the model updates will progressively contract because the model is meant to converge with training, meaning that the quantization level should decrease as training progresses.
@@
AdaQuantFL [Jhunjhunwala et al.(2021)] X X An adaptive quantization technique modifies the number of quantization levels throughout training with the goal of achieving both a low error floor and effective communication.
@@
FedQOGD [Park et al.(2022)] X X Gradient descent quantized communication efficient FL framework using partial node participation and stochastic quantization. FedQOGD has a minimal communication overhead and maintains edge-node privacy while yielding the same asymptotic performance as its centralized equivalent (i.e., all local data are gathered at the central server).
@@
In order to alleviate the communication delay constraint in federated learning, gradient compression—which includes gradient quantization and sparsification—has been extensively studied. Quantization modifies a floating-point number of bits to compress parameters [Tonellotto et al.(2021)]. Low-precision representation of deep neural networks (DNNs) is a potential method to allow for effective memory reduction and acceleration. Hence, the reduction in energy consumption.
Energy and Carbon Footprint Study in Federated Learning